Illustrate the graph of the given one-to-one


1. f(x)=3x-4 2. f(x) = 5x+3


3. f(x)=7x-5


4. f(x)=1-2


5. f(x)=123


6. f(x)=x+5


1
Expert's answer
2021-10-24T17:45:32-0400

1. "f(x)=3x-4"




2. "f(x) = 5x+3"





3."f(x)=7x-5"




4. "f(x)=1-2"


"f(x)=-1"

The function "f(x)=1-2" is not a one-to-one function



"f(0)=-1=f(1), 0\not=1"


5. "f(x)=123"


The function "f(x)=123" is not a one-to-one function



"f(0)=123=f(1), 0\not=1"


6. "f(x)=x+5"